# Signal Recovery From Product of Two Vandermonde Matrices

### 27 Jan 2024 v1

@@ 3. Output: $\bm{\theta}^{+},e^{i\alpha}\bm{g}^{+}$ for some unknown $\alpha$
For the given $\bm{\theta}^{+}$, $\bm{g}^{+}$, there clearly exists a vector $\hat{\bm{z}}=[\hat{z}_{0},\ldots,\hat{z}_{8s-4}]$ of $38s-$ 3 samples that satisfies the assumptions of Theorem 8. Thus, $\bm{G}(\hat{\bm{z}})$ in (III-A) has a one-dimensional null space. Since $\bm{G}(\hat{\bm{z}})$ is a $(8s-3)\times(8s-2)$ matrix, a one-dimensional null space of $\bm{G}(\hat{\bm{z}})$ implies that there is a $(8s-3)\times(8s-3)$ sub-matrix $\bm{M}(\hat{\bm{z}})$ of $\bm{G}(\hat{\bm{z}})$ with full rank $38s-$ 3. $\bm{M}(\hat{\bm{z}})$ is obtained from $\bm{G}(\hat{\bm{z}})$ by removing one column from $\bm{G}(\hat{\bm{z}})$. We will show that we can replace $\hat{\bm{z}}$ with the arbitrary $\bm{z}$ and obtain that $\bm{M}(\bm{z})$ is full rank w.p.1.
From (59), it follows that $y_{j}^{+}$ is a ratio of two Laurent polynomials in the variable $z_{j}$. Thus, the determinant $\left|\bm{M}(\bm{z})\right|$ is a ratio $P(\bm{z})/Q(\bm{z})$ of multivariate Laurent polynomials $P(\bm{z})$ and $Q(\bm{z})$. Since $\bm{M}(\hat{\bm{z}})$ is invertible, $P(\hat{\bm{z}})\neq 0$, so $P(\bm{z})$ is not zero. The set $\mathbb{T}^{8s-3}$ is homeomorphic to the $(8s-3)$-dimensional cube $\mathbb{U}^{8s-3}$, where $\mathbb{U}=[-1,1)$ is the unit interval. Hence, the zero set $\mathcal{S}=\{\bm{z}|\bm{z}\in\mathbb{T}^{8s-3},P(\bm{z})=0\}$ is either empty or is a set of points in $\mathbb{U}^{8s-3}$ taken from a hypersurface of dimension at most $48s-$ 4. An arbitrary $\bm{z}\in\mathbb{T}^{8s-3}$, corresponding to an arbitrary point in $\mathbb{U}^{8s-3}$, is inside $\mathcal{S}$ w.p.0, proving the theorem. ∎
@@
Theorem 9 is true for arbitrary $\bm{z}\in\mathbb{T}^{m}$ with $m\geq 8s-3$ as well, since any subset of $38s-$ 3 samples from $\bm{z}$ ($38s-$ 3 rows from $\bm{G}(\bm{z})$) satisfies Theorem 9 w.p.1. Similar to Proposition 5 we have
Proposition 6.
@@
Given the assumptions in Theorem 9, $|\bm{g}^{+}|$ can be found, up to a real scalar $c$, with $O(\mathrm{poly}(s))$ complexity.
@@
From Theorem 9 we obtain $\bm{\theta}^{+}$ and $L(z)=cL^{+}(z)$. Using $(\theta_{k}^{+})^{-1}=\overline{\theta_{k}^{+}}$ and
@@
Since the operations leading to $c|\bm{g}^{+}|$ are of $O(\mathrm{poly}(s))$ complexity, it follows that $c|\bm{g}^{+}|$ can be found with $O(\mathrm{poly}(s))$ complexity. ∎
@@
We have the option to choose sign in (119), resulting in the opposite sign in (118). First, assume that -- is chosen in (119). Denote the roots of the obtained $|\tilde{u}^{+}(z)|^{2}$, which appear as conjugate pairs, as $(r_{0},1/\overline{r_{0}}),\ldots,(r_{s-2},1/\overline{r_{s-2}})$. For arbitrary $\bm{g}^{+}$, since $\tilde{L}^{+}(z)=\hat{u}^{+}(z)\overline{\tilde{u}^{+}(z)}$ and due to Lemma 2 and Theorem 1, $s-1$ roots of $|\tilde{u}^{+}(z)|^{2}$ appear as roots of $\tilde{L}^{+}(z)$ w.p.1. Assume WLOG that the roots $r_{0},\ldots,r_{s-2}$ appear as roots of $\tilde{L}^{+}(z)$. Hence, we construct $\tilde{u}^{+}(z)$ (up to a complex scalar) as $\hat{u}^{+}(z)=\prod_{j=0}^{s-2}(z-r_{j})$ (the other $s-1$ roots of $\tilde{L}^{+}(z)$ produce $\hat{u}^{+}(z)$). Solving the equations
@@
An important special case is when $L^{2}(z)=4K(z)$, which implies that $|\hat{u}(z)|^{2}=|\tilde{u}(z)|^{2}=L(z)/2$. Theorem 1 shows that this occurs for arbitrary $\bm{g}$ if and only if $\theta_{k}^{n}$, $10\leq k\leq s-$ 1, are all equal. Since $\bm{\theta}\in\mathbb{T}^{s}$, this means that $\theta_{k},0\leq k\leq s-1$, are shifted harmonics and thus $\bm{x}$ is $s$-sparse in a DFT basis. In this case, the situation is the same as in R5-Alg-Harmonic since there are $2^{s-1}$ different ways to choose the roots for $\tilde{u}(z)$ (each one of them producing a solution to the equations after normalization); thus, the recovery complexity suddenly becomes exponential in $s$. Hence, the implication of this observation is that when $\bm{x}$ is $s$-sparse in a DFT basis, the recovery complexity of $\bm{x}$ is $O(\mathrm{poly}(s)2^{s})$ (as mentioned in 2.b in Section I). Nevertheless, due to Theorem 1, as soon as not all $\theta_{j}^{n}$, $10\leq j\leq s-$ 1, are equal (i.e., at least one frequency is not a shifted harmonic), $L^{2}(z)\neq 4K(z)$ w.p.1 since then the roots of $\tilde{u}(z)$ and $\hat{u}(z)$ are different w.p.1.
If $S$ of the $g_{0},\ldots,g_{s-1}$ are non-zero, it is readily seen that $\bm{w}^{+}$ in (71) belongs to the null space of $\bm{G}(\bm{z})$ for $s\geq S$; thus the same iterative procedure as in step 2 of R5-Alg-Harmonic can be used to find $S$. Hence, we have the following algorithm that recovers $\bm{g}^{+},\bm{\theta}^{+}$
@@ 2. Output: $\bm{\theta}^{+},e^{i\alpha}\bm{g}^{+}$ for some unknown $\alpha$
    Construct the $m\times(8s-1)$ matrix $\bm{G}(\bm{z})$ in (III-A).
@@
    Construct the Laurent polynomials $\hat{L}(z)=\sum_{k=0}^{2S}w_{k}z^{k-S},\tilde{L}(z)=\sum_{k=0}^{2S-2}w_{k+2S+1}z^{k-S+1},L(z)=\sum_{k=0}^{2S-2}w_{k+4S}z^{k-S+1}$, respectively. The conjugate of the unique roots of $\hat{L}(z)$ equal $\bm{\theta}^{+}$. $|\bm{g}^{+}|$ can now be computed (up to a global real scalar) from Proposition 6.
@@
      $L^{2}(z)\neq 4K(z)$ ($(\theta_{j}^{+})^{n}$, $10\leq j\leq S-$ 1, are not all equal):
@@
      $L^{2}(z)=4K(z)$ ($(\theta_{j}^{+})^{n}$, $10\leq j\leq S-$ 1, are all equal):
      • In this case, $L(z)/2=|\hat{u}(z)|^{2}=|\tilde{u}(z)|^{2}$. The roots of $L(z)/2$ can be paired as $(l_{0},1/\overline{l_{0}}),\ldots,(l_{S-2},1/\overline{l_{S-2}})$. From each pair, choose one element. This choice can be done in $2^{S-1}$ different ways and let $\bm{l}^{k}$, $10\leq k\leq 2^{S-1}-$ 1, denote the $k$:th choice.
      • For each $\bm{l}^{k}$, $10\leq k\leq 2^{S-1}-$ 1, solve for $\bm{g}$ the $S-1$ polynomial equations
        $\sum_{k=0}^{S-1}g_{k}t_{k}(l_{j})=0,\quad 0\leq j\leq S-2.$
@@
    Let $\bm{g}^{+}=\arg\min_{\bm{g}\in\mathcal{G}}|y_{m}-|\bm{a}^{T}\bm{V}(\bm{\theta}^{+})\bm{g}|^{2}$. Output $\bm{g}^{+}$ and $\bm{\theta}^{+}$.
@@ 3. Output: $\bm{x}$ (up to a global phase)
R3-Alg shows that there is a unique $\bm{x}$ (up to a global phase) satisfying the constraint in (4) for $\bm{A}=[\bm{V}(\bm{z})^{T}\bm{V}(\bm{\theta});\bm{a}^{T}]$ with $m=O(s)$. This observation is what solves, and significantly improves upon, open problem 3.5 in [18].

### 06 Feb 2024 v2

@@ 3. Output: $\bm{\theta}^{+},e^{i\alpha}\bm{g}^{+}$ for some unknown $\alpha$
Herein, we will denote $\bm{G}(\bm{z})$ in (III-A) as $\bm{G}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ to reflects its dependency on $\bm{\theta}^{+}$, $\bm{g}^{+}$ as well. A one-dimensional null space of $\bm{G}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ implies that there exists a $(8s-3)\times(8s-3)$ sub-matrix in $\bm{G}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ of full rank $38s-$ 3; this sub-matrix is obtained by removing a column from $\bm{G}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$. Let $d_{k}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ denote the determinant of the $(8s-3)\times(8s-3)$ sub-matrix obtained after removing column $k$, $21\leq k\leq 8s-$ 2, from $\bm{G}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$. Hence, the statement that $\bm{G}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ has a null space of at least two dimensions is equivalent to $d_{k}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})=0,1\leq k\leq 8s-2$. From (59), it follows that $y_{j}^{+}$, and thus $d_{k}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ as well, is a ratio of two multivariate Laurent polynomials in the variables $\bm{z},\bm{\theta}^{+},\bm{g}^{+}$.
First, we show that $\bm{G}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ has a one-dimensional null space for almost all $\bm{z}\in\mathbb{T}^{8s-3}$ given a certain realization $\bm{\theta}^{+}=\hat{\bm{\theta}}^{+}$ and $\bm{g}^{+}=\hat{\bm{g}}^{+}$ (with $s$ non-zero elements). Theorem 8 shows that for almost all $\bm{z}$ that contain shifted harmonics (with different shifts) as specified in the theorem, $\bm{G}(\bm{z},\hat{\bm{\theta}}^{+},\hat{\bm{g}}^{+})$ has a one dimensional null space; let $\bm{z}=\hat{\bm{z}}$ be one such vector. Hence, $d_{j}(\hat{\bm{z}},\hat{\bm{\theta}}^{+},\hat{\bm{g}}^{+})\neq 0$ for some $j$, $21\leq j\leq 8s-$ 2, implying that $d_{j}(\bm{z},\hat{\bm{\theta}}^{+},\hat{\bm{g}}^{+})$ is not zero for all $\bm{z}\in\mathbb{T}^{8s-3}$. Consider the zero set $\mathcal{S}_{j}=\{\bm{z}|\bm{z}\in\mathbb{T}^{8s-3},d_{j}(\bm{z},\hat{\bm{\theta}}^{+},\hat{\bm{g}}^{+})=0\}$. Since $\mathbb{T}^{8s-3}$ is homeomorphic to the $38s-$ 3 dimensional cube $\mathbb{U}^{8s-3}$ (with the obvious homeomorphism $e^{i\pi\alpha}\in\mathbb{T},\alpha\in\mathbb{U}=[-1,1)$), $\mathcal{S}_{j}$ is either empty or a set of points in $\mathbb{T}^{8s-3}$ corresponding to points from a hypersurface of dimension $48s-$ 4 in the ambient space $\mathbb{U}^{8s-3}$. In either case, $\mathcal{S}_{j}$ has measure zero, implying that for almost all $\bm{z}\in\mathbb{T}^{8s-3}$, $d_{j}(\bm{z},\hat{\bm{\theta}}^{+},\hat{\bm{g}}^{+})\neq 0$ and thus $\bm{G}(\bm{z},\hat{\bm{\theta}}^{+},\hat{\bm{g}}^{+})$ has a one-dimensional null space. Hence, for any given $\bm{\theta}^{+}$ and $\bm{g}^{+}$ (with $s$ non-zero elements), $\bm{G}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ has a one-dimensional null space for almost all $\bm{z}\in\mathbb{T}^{8s-3}$.
We now show that for almost any given $\bm{z}\in\mathbb{T}^{8s-3}$, $\bm{G}(\bm{z},\bm{\theta}^{+},\bm{g}^{+})$ has a one-dimensional null space for almost all $\bm{\theta}^{+}\in\mathbb{T}^{s}$ and $\bm{g}^{+}\in\mathbb{C}^{s}$ (with $s$ non-zero elements). Since almost all $\bm{z}\in\mathbb{T}^{8s-3}$ do not belong to the set $\mathcal{S}_{j}$ above, choose any such element $\tilde{\bm{z}}\not\in\mathcal{S}_{j}$. Hence, $d_{j}(\tilde{\bm{z}},\hat{\bm{\theta}}^{+},\hat{\bm{g}}^{+})\neq 0$, implying that $d_{j}(\tilde{\bm{z}},\bm{\theta}^{+},\bm{g}^{+})$ is not zero for all $\bm{\theta}^{+}\in\mathbb{T}^{s}$ and $\bm{g}^{+}\in\mathbb{C}^{s}$. Consider the set $\mathcal{T}_{j}=\{\bm{\theta}^{+},\bm{g}^{+}|\bm{\theta}^{+}\in\mathbb{T}^{s},\bm{g}^{+}\in\mathbb{C}^{s},d_{j}(\tilde{\bm{z}},\bm{\theta}^{+},\bm{g}^{+})=0\}$. This is a subset of the set $\mathbb{T}^{8s-3}\times\mathbb{C}^{s}$, where $\times$ here denotes the Cartesian product between sets. From above, it follows that $\mathbb{T}^{8s-3}\times\mathbb{C}^{s}$ is homeomorphic to the set $\mathbb{U}^{8s-3}\times\mathbb{R}^{2s}$ of (real) dimension $38s-3+2s=10s-$ 3 (the complex numbers $\mathbb{C}$ are represented by $\mathbb{R}^{2}=\mathbb{R}\times\mathbb{R}$). Hence, $\mathcal{T}_{j}$ is either empty or a set of points in $\mathbb{T}^{8s-3}\times\mathbb{C}^{s}$ corresponding to points from a hypersurface of dimension $58s-3+2(s-1)=10s-$ 5 in the ambient space $\mathbb{U}^{8s-3}\times\mathbb{R}^{2s}$. In both cases, it has measure zero in $\mathbb{U}^{8s-3}\times\mathbb{R}^{2s}$, implying that $\bm{G}(\tilde{\bm{z}},\bm{\theta}^{+},\bm{g}^{+})$ has a one-dimensional null space for almost all $\bm{\theta}^{+},\bm{g}^{+}$. ∎
@@
Theorem 9 is true for almost all $\bm{z}\in\mathbb{T}^{m}$ with $m\geq 8s-3$ as well, since almost all subsets of $38s-$ 3 samples from $\bm{z}$ ($38s-$ 3 rows from $\bm{G}(\bm{z})$) satisfy Theorem 9.
Similar to Corollary 3 we have
Corollary 4.
@@
Given the realizations $\bm{\theta}^{+},\bm{g}^{+}$ of $\bm{\theta},\bm{g}$, respectively. For almost all vectors $\bm{z}\in\mathbb{T}^{m}$ with $m\geq 8s-3$, $\bm{\theta}^{+}$ and $|\bm{g}^{+}|$ can be found, up to a real positive scalar $c$, with $O(\mathrm{poly}(s))$ complexity.
@@
Theorem 9 shows that for almost all $\bm{z}$, one obtains $\bm{\theta}^{+}$ and $L(z)=cL^{+}(z)$. Using $(\theta_{k}^{+})^{-1}=\overline{\theta_{k}^{+}}$ and
@@
Since the operations leading to $c|\bm{g}^{+}|$ are of $O(\mathrm{poly}(s))$ complexity, it follows that $\sqrt{c}|\bm{g}^{+}|$ can be found with $O(\mathrm{poly}(s))$ complexity. ∎
@@
We have the option to choose sign in (119), resulting in the opposite sign in (118). First, assume that -- is chosen in (119). Denote the roots of the obtained $|\tilde{u}^{+}(z)|^{2}$ (which are distinct according to Lemma 2 and appear as conjugate pairs) as $(r_{0},1/\overline{r_{0}}),\ldots,(r_{s-2},1/\overline{r_{s-2}})$. Since $\tilde{L}^{+}(z)=\hat{u}^{+}(z)\overline{\tilde{u}^{+}(z)}$ and due to Theorem 1, $s-1$ roots of $|\tilde{u}^{+}(z)|^{2}$ appear as roots of $\tilde{L}^{+}(z)$ for almost all $\bm{g}^{+}$. Assume WLOG that the roots $r_{0},\ldots,r_{s-2}$ appear as roots of $\tilde{L}^{+}(z)$. Hence, we construct $\tilde{u}^{+}(z)$ (up to a complex scalar) as $\hat{u}^{+}(z)=\prod_{j=0}^{s-2}(z-r_{j})$ (the other $s-1$ roots of $\tilde{L}^{+}(z)$ produce $\hat{u}^{+}(z)$). Solving the equations
@@
An important special case is when $(L^{+}(z))^{2}=4K^{+}(z)$, which implies that $|\hat{u}^{+}(z)|^{2}=|\tilde{u}^{+}(z)|^{2}=L^{+}(z)/2$. Theorem 1 shows that for his to occur with a non-zero probability, $(\theta_{k}^{+})^{n}$, $10\leq k\leq s-$ 1, must all be equal. Since $\bm{\theta}^{+}\in\mathbb{T}^{s}$, this means that $\theta_{k}^{+},0\leq k\leq s-1$, are shifted harmonics and thus $\bm{x}^{+}$ is $s$-sparse in a DFT basis. In this case, the situation is the same as in R5-Alg-Harmonic since there are $2^{s-1}$ different ways to choose the roots for $\tilde{u}^{+}(z)$ (each one of them producing a solution to the equations after normalization); thus, the recovery complexity suddenly becomes exponential in $s$. Hence, the implication of this observation is that when $\bm{x}^{+}$ is $s$-sparse in a DFT basis, the recovery complexity of $\bm{x}^{+}$ is $O(\mathrm{poly}(s)2^{s})$ (as mentioned in 2.b in Section I). Nevertheless, due to Theorem 1, as soon as not all $(\theta_{j}^{+})^{n}$, $10\leq j\leq s-$ 1, are equal (i.e., at least one frequency is not a shifted harmonic), $(L^{+}(z))^{2}\neq 4K^{+}(z)$ with probability 1 since then the roots of $\tilde{u}^{+}(z)$ and $\hat{u}^{+}(z)$ are different for almost all $\bm{g}^{+}$.
If $S$ of the $g_{0}^{+},\ldots,g_{s-1}^{+}$ are non-zero, it is readily seen that $\bm{w}^{+}$ in (71) belongs to the null space of $\bm{G}(\bm{z})$ for $s\geq S$; thus the same iterative procedure as in step 2 of R5-Alg-Harmonic can be used to find $S$. Hence, we have the following algorithm that recovers $\bm{g}^{+},\bm{\theta}^{+}$
@@ 2. Output: $\bm{\theta}^{+},e^{i\alpha}\bm{g}^{+}$ for some unknown $\alpha$
    Construct the $m\times(8s-2)$ matrix $\bm{G}(\bm{z})$ in (III-A).
@@
    Construct the Laurent polynomials $\hat{L}(z)=\sum_{k=0}^{2S}w_{k}z^{k-S},\tilde{L}(z)=\sum_{k=0}^{2S-2}w_{k+2S+1}z^{k-S+1},L(z)=\sum_{k=0}^{2S-2}w_{k+4S}z^{k-S+1}$, respectively. The conjugate of the unique roots of $\hat{L}(z)$ equal $\bm{\theta}^{+}$. $|\bm{g}^{+}|$ can now be computed (up to a global real positive scalar) from Corollary 4.
@@
      $L^{2}(z)\neq 4K(z)$ (corresponding to $(\theta_{j}^{+})^{n}$, $10\leq j\leq S-$ 1, not all being equal):
@@
      $L^{2}(z)=4K(z)$ (corresponding to $(\theta_{j}^{+})^{n}$, $10\leq j\leq S-$ 1, all being equal):
      • In this case, $L(z)/2=|\hat{u}(z)|^{2}=|\tilde{u}(z)|^{2}$. The roots of $L(z)/2$ can be paired as $(l_{0},1/\overline{l_{0}}),\ldots,(l_{S-2},1/\overline{l_{S-2}})$. From each pair, choose one element. This choice can be done in $2^{S-1}$ different ways and let $\bm{q}^{k}$, $10\leq k\leq 2^{S-1}-$ 1, where $q_{j}^{k}=l_{j}$ or $q_{j}^{k}=1/\overline{l_{j}}$ for $20\leq j\leq S-$ 2, denote the $k$:th choice.
      • For each $\bm{q}^{k}$, $10\leq k\leq 2^{S-1}-$ 1, solve for $\bm{g}$ the $S-1$ polynomial equations
        $\sum_{n=0}^{S-1}g_{n}t_{n}(q_{j}^{k})=0,\quad 0\leq j\leq S-2.$
@@
    Let $\bm{g}^{+}=\arg\min_{\bm{g}\in\mathcal{G}}|y_{m}-|\bm{a}^{T}\bm{V}(\bm{\theta}^{+})\bm{g}|^{2}|$. Output $\bm{g}^{+}$ and $\bm{\theta}^{+}$.